OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 The solution to the system of equations is A.(13/34, -15/34) B.(3/22, -7/22) C.(13/22, 1/22)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Can you please help me?

OpenStudy (ivettef365):

you can start by multiplying the first equation by (4) and the second equation by (-3)

OpenStudy (ivettef365):

(4) 7x -3y = 4 then you have 28x -12y = 16 (-3) 2x -4y = 1 then you have -6x +12y = -3 now subtract ---------------- what do you get: ????????

OpenStudy (anonymous):

So when im about to Subtract i. do the left 28x-12y=16 & -6x=12y+-3